OpenStudy (anonymous):

Every day, Marco runs for 45 minutes and covers 5 miles. How long would it take him to run 8 miles? 9 min 18 min 48 min 72 min

OpenStudy (saifoo.khan):

Duration : Miles 45 : 5 x : 8 Cross multiply.

Directrix (directrix):

72 min ------- 45 minutes is to 5 miles as M minutes is to 8 miles 45/5 = M/8 9/1 = M/8 M = 9*8 M =72 minutes